A farmer raises wheat and corn on 430 acres of land. He wants to plant 62 more acres of wheat than corn. How many acres of each should he plant?

A. Write a system of equations for the problem.
B. Solve your system of equations. Make sure to show/explain all of your work for credit

Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

answered yesterday

Answer by Theo(13342)

let x equal the number of acres of wheat.
let y equal the number of acres of corn.

he wants to plant 62 more acres of wheat than corn.

that means that x = y + 62.

you have 2 equations.

the first equation is x + y = 430.

this means that the number of acres of wheat plus the number of acres of corn must be equal to 430 acres in total.

the second equation is x = y + 62.

this means that the number of acres of wheat must be 62 more than the number of acres of corn.

in the equation of x + y = 430, replace x with it's equivalent value of y + 62 to get:

x + y = 430 becomes y + 62 + y = 430.

combine like terms to get 2y + 62 = 430.

subtract 62 from both sides of this equation to get 2y = 368.

divide both sides of this equation by 2 to get y = 184.

since x is equal to y + 62, this means that x must be equal to 184 + 62 which is equal to 246.

you have x = 246 acres and y = 184 acres.

x + y = 430 becomes 264 + 184 = 430 which becomes 430 = 430.

this confirms that the value of x and y are good in that equation.

you also get x = y + 62 becomes 246 = 184 + 62 which becomes 246 = 246.

this confirms that the value of x and y are good in that equation.

the values of x and y are good in both equations, so the solution is good.

your solution is that the number of acres of wheat is equal to 246 and the number of acres of corn is 184.